media%2F47b%2F47b34e70-36d3-4b60-8fd6-df
Answer #1

C1 C2 C3 C4Given data

C_{1}=9nF

C_{2}=4nF

C_{3}=2nF

C_{4}=3nF

V_{0}=9volt

Equivalent capacitance between point C and D can be given by

C_{CD}=C_{3}+C_{4}=(2+3)nF=5nF

Equivalent capacitance between A and B can be given by

C_{AB}=C_{CD}+C_{2}=(5+4)nF=9nF

Equivalent capacitance of the total circuit can be given by

C=\frac{C_{AB}\times C_{1}}{C_{AB}+C_{1}}=\frac{9\times 9}{9+9}nF=4.5nF

Total charge can be given by

Q=CV_{0}=(4.5\times 10^{-9}\times 9)C=40.5\times 10^{-9}C

Voltage accross A and B can be given by

V_{AB}=\frac{Q}{C_{AB}}=\frac{40.5\times 10^{-9}}{9\times 10^{-9}}volt=4.5volt

Voltage accross C_{4} can be given by

V_{C_{4}}=4.5volt